# STRRINGS: STReams in Residual Images of Nearby GalaxieS

### v1

@@ Acknowledgements
We thank Kyle Adams and Maria Skryabina for sharing their SB measurements of streams from their paper Skryabina et al. (2024).
E.S is grateful to the Leverhulme Trust for funding under the grant number RPG-2021-205. OM is grateful to the Swiss National Science Foundation for financial support under the grant number PZ00P2_202104. AAA acknowledges support from the Herchel Smith Fellowship at the University of Cambridge and a Fitzwilliam College research fellowship supported by the Isaac Newton Trust. EYD thanks the Science and Technology Facilities Council (STFC) for a PhD studentship (UKRI grant number 2605433). HZ thanks the Science and Technology Facilities Council (STFC) for a PhD studentship. DC acknowledges funding from the Harding Distinguished Postgraduate Scholars Program. JLM acknowledges the support of a fellowship from the “la Caixa” Foundation (ID 100010434), the fellowship code is LCF/BQ/EU24/12060091.
@@ Appendix C Photometry from original and residual images
We investigate the differences between the photometry of the stream estimated from the original images using a standard method versus from the residual images. For the original images, we download the DECaLS images in bands at a pixel size of 0.262 from the DESI Legacy survey files151515DESI Legacy survey files. For each of our 35 galaxies, we retrieve all the images in a field-of-view large enough around the galaxy, combine them using SWarp and crop the result to be centred on the object. We then perform source extraction and masking using SEP (Barbary, 2016), the Python implementation of SExtractor (Bertin & Arnouts, 1996), with a detection threshold of 2.5 and a background mesh size of 128128 pixels. We then apply the stream segmentation mask from Jafar. We retrieve the values of the flux, surface brightness and colours inside the mask. We compare these values to the ones obtained directly from the residual images (as described in Section 3.2).

### v2

@@ Acknowledgements
We thank the referee, Garreth Martin, for constructive comments that helped improve the manuscript. We thank Kyle Adams and Maria Skryabina for sharing their SB measurements of streams from their paper (Skryabina et al., 2024). E.S. and V.B. are grateful for the support from the Leverhulme Trust Research Project Grant RPG-2021-205 “The Faint Universe Made Visible with Machine Learning”. OM is grateful to the Swiss National Science Foundation for financial support under the grant number PZ00P2_202104. AAA acknowledges support from the Herchel Smith Fellowship at the University of Cambridge and a Fitzwilliam College research fellowship supported by the Isaac Newton Trust. EYD thanks the Science and Technology Facilities Council (STFC) for a PhD studentship (UKRI grant number 2605433). HZ thanks the Science and Technology Facilities Council (STFC) for a PhD studentship. DC acknowledges funding from the Harding Distinguished Postgraduate Scholars Program. JLM acknowledges the support of a fellowship from the “la Caixa” Foundation (ID 100010434), the fellowship code is LCF/BQ/EU24/12060091.
@@ Appendix C Photometry from original and residual images
We investigate the differences between the photometry of the stream estimated from the original images using a standard method versus from the residual images. For the original images, we download the DECaLS images in bands at a pixel size of 0.262 from the DESI Legacy survey files171717DESI Legacy survey files. For each of our 35 galaxies, we retrieve all the images in a field-of-view large enough around the galaxy, combine them using SWarp and crop the result to be centred on the object. We then perform source extraction and masking using SEP (Barbary, 2016), the Python implementation of SExtractor (Bertin & Arnouts, 1996), with a detection threshold of 2.5 and a background mesh size of 128128 pixels. We then apply the stream segmentation mask from Jafar. We retrieve the values of the flux, surface brightness and colours inside the mask. We compare these values to the ones obtained directly from the residual images (as described in Section 3.2).